OpenStudy (anonymous):

please help me find all exaclty solutions for these 2 equation: 9^x + 14 = 3^(x+2) (logx)^3 = 3 logx

OpenStudy (anonymous):

the first equation, do logs log both sides, see where that gets you

OpenStudy (tkhunny):

Is that \[9^{x+14}\] or \[9^{x} + 14\]?

OpenStudy (anonymous):

the upper one

OpenStudy (tkhunny):

Well, that makes it possible. If it were the other, there woudl be no exact solution. Observe that 9 = 3^2. This nearly completes the problem.

OpenStudy (anonymous):

should i log both side? cuz if i do that, log 14 is a big mess

OpenStudy (tkhunny):

Did you use the hint?\[9^{x+14} = 3^{x+2}\] Using the hint, we have \[\left(3^{2}\right)^{x+14} = 3^{2x+28} = 3^{x+2}\] It is now pretty easy IF you see it. Do you?